[Ian, Daniel, Jeff, Torrey]
Moving of the optics sleds has been completed without any hiccups. Particle count pre moving was 0/0/0 and 20 minutes after things have settled 0/0/0. The mylar is still over the cavities, we will unwrap when we plan on turning on lasers. Pictures attached.
Please note that the fans in the mobil cleans rooms in B111 should never be turned off now.
Looks great. Can you remind me what the units are on the particle counter? What are the three particle sizes, and is it calibrated in terms of particles per volume?
x/x/x is the 0.3 um, 0.5 um, and 1.0 um cumulative count over 1 minute.